OpenStudy (anonymous):

Please help? George is comparing three investment accounts offering different rates. Account A: APR of 3.75 compounding monthly Account B APR of 3.85 compounding quarterly Account C APR of 3.95 compounding daily Which account will give George at least a 4% annual yield?

OpenStudy (anonymous):

you need to compute three numbers \[(1+\frac{.0375}{4})^4\] \[(1+\frac{.0385}{12})^{12}\] \[(1+\frac{.0395}{365})^{365}\]

OpenStudy (anonymous):

so which ever equals at least "4" will give me the answer then?

OpenStudy (anonymous):

each will look like \(1+.0somthing\) choose the one or ones with the something greater than or equal to 4

OpenStudy (anonymous):

none will give you 4 if the answer looks like \(1.04...\) then that is a yield of larger than \(4\%\) if the answer is \(1.03...\) then no

OpenStudy (anonymous):

wait the first two equations are mixed up am i right?

OpenStudy (anonymous):

oh yeah my fault, sorry

OpenStudy (anonymous):

\[(1+\frac{.0375}{12})^{12}\]

OpenStudy (anonymous):

it's quite alright. I know how to set up the problem now. but about the yeild, the number won't equal 4, but if the last digit in the number is greater than or equal to 4 than that is the answer i would choose?

OpenStudy (anonymous):

\[(1+\frac{.0385}{4})^{4}\] good catch

OpenStudy (anonymous):

lets compute one

OpenStudy (anonymous):

here is the first one http://www.wolframalpha.com/input/?i=%281%2B.0375%2F12%29^12

OpenStudy (anonymous):

you can see that the answer is \[1.033815...\]

OpenStudy (anonymous):

so that one can be ruled out

OpenStudy (anonymous):

so this one is out since it is smaller than \(1.04\)

OpenStudy (anonymous):

right

OpenStudy (anonymous):

I'll try the second one..

OpenStudy (anonymous):

ok you can edit the one i did to compute it if you like

OpenStudy (anonymous):

is the little 4 on the outside an exponent?

OpenStudy (anonymous):

yes, that needs to be a 4, and also the denominator under \(.0385\) should be a 4 also

OpenStudy (anonymous):

1.039 is what it gave me

OpenStudy (anonymous):

me too

OpenStudy (anonymous):

it is the last one, it gave me 1.04

OpenStudy (anonymous):

then i guess that is the one you want

OpenStudy (anonymous):

thanks for the help

OpenStudy (anonymous):

you good with this?

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

yes indeed :3